You are a manager for Tesla and recently hired an economist to work with engineering and operations experts to estimate the production function for lithium battery production, and the production function is Q = 5 K 3/4 L 1/4. Tesla has already spent $12,000 on the 10 units of capital it owns and cannot acquire more capital in the short run; if workers are paid $ 1,000 per day and batteries can be sold for $10,000 each:

a) What is the profit maximizing level of the output and labor usage? b) What is the maximum profit?
Answer #1

We have a production function Q = 5 K^(3/4) L^(1/4)

Here K = 10 and is fixed. Hence Q = 5*10^(3/4) L^(1/4) or Q = 28.117L^(1/4)

This gives L* = (Q/28.117)^4

Cost function is C = wL + rK

C = 1000*(Q/28.117)^4 + 12000

C = 0.0016Q^4 + 12000

Profit = revenue - cost

= 10000Q - 0.0016Q^4 - 12000

Profit is maximum when its derivative is 0

10000 - 0.0016*4Q^3 = 0

This gives Q = 116 units and L = 290 units

Profit = $858298